OpenStudy (rainbow_rocks03):

PLZ HELP I WILL MEDAL YOU The circle graph below shows the percentages of Mrs. Frederick's monthly budget for July. Her total monthly income is $2,800. In August's monthly budget, Mrs. Frederick spent $126 more on clothes than she did in July. In August, what percentage of her monthly income of $2,800 was spent on clothes? A. 15% B. 12.5% C. 11.5% D. 10%

OpenStudy (rainbow_rocks03):



OpenStudy (rainbow_rocks03):

@whpalmer4 @rishavraj @ikram002p @Astrophysics

OpenStudy (whpalmer4):

how much did she spend on clothes in July?

OpenStudy (rainbow_rocks03):

$224 @whpalmer4 ?

OpenStudy (whpalmer4):

ok, if in August she spends %126 more than she spent in July, how much did she spend in August?

OpenStudy (rainbow_rocks03):

ummm

OpenStudy (rainbow_rocks03):

I dont know. 😞@whpalmer4

OpenStudy (whpalmer4):

what is 126 more than 224?

OpenStudy (rainbow_rocks03):

126 more than 224?

OpenStudy (rainbow_rocks03):

Would I subtract 224 - 126 ?

OpenStudy (whpalmer4):

if you have 10 dollars, and I have 5 dollars more than you do, how many dollars do I have?

OpenStudy (rainbow_rocks03):

15?

OpenStudy (whpalmer4):

yes. so what is 126 more than 224?

OpenStudy (rainbow_rocks03):

98

OpenStudy (whpalmer4):

no, how do you come up with that? is 98 more than 224?

OpenStudy (rainbow_rocks03):

ohhh

OpenStudy (rainbow_rocks03):

uh, im getting a little confused. sorry :(

OpenStudy (whpalmer4):

if I say some number is 35 more than some other number, that means the first number equals 35 plus the other number.

OpenStudy (rainbow_rocks03):

ok

OpenStudy (whpalmer4):

so, once again, if she spends 126 more than 224, how much does she spend?

OpenStudy (whpalmer4):

the amount she spends is 126 more than 224

OpenStudy (rainbow_rocks03):

350?

OpenStudy (whpalmer4):

yes!!! 😀

OpenStudy (rainbow_rocks03):

Im sorry, im trying my best.. its just i have been working sense 11.00 am

OpenStudy (whpalmer4):

now, what percentage is 350 of her monthly budget of 2800?

OpenStudy (rainbow_rocks03):

ok I was thinking 350 but I was like umm im not sure

OpenStudy (rainbow_rocks03):

clothes

OpenStudy (whpalmer4):

what percentage is 350 of 2800? not which category are we discussing...

OpenStudy (rainbow_rocks03):

8%

OpenStudy (whpalmer4):

no, 224 was 8% of 2800

OpenStudy (rainbow_rocks03):

oh wait.

OpenStudy (rainbow_rocks03):

I didnt mean to put 8%

OpenStudy (whpalmer4):

\[\frac{350}{2800}=\frac{x}{100}\] \(x\)% will be your answer

OpenStudy (rainbow_rocks03):

i cross multiply?

OpenStudy (whpalmer4):

that is one way you could do it

OpenStudy (rainbow_rocks03):

so 350*100=35000

OpenStudy (rainbow_rocks03):

and 2800*x=2800x

OpenStudy (whpalmer4):

yes, and those two quantities equal each other. how much is \(x\)?

OpenStudy (rainbow_rocks03):

2800?

OpenStudy (whpalmer4):

\[35000=2800x\] does it look like \( x=2800\) will be the solution? if you say yes, it means you think that \(2800*2800=35000\)

OpenStudy (rainbow_rocks03):

ohh...2800 * 2800 = 7,840,000

OpenStudy (whpalmer4):

which does not equal 35000 😀

OpenStudy (whpalmer4):

why don't you divide each side by some number...

OpenStudy (rainbow_rocks03):

I divide each side of 7,840,000 by some number

OpenStudy (rainbow_rocks03):

Im thinking really hard about this. 😓

OpenStudy (whpalmer4):

go back to my equation. \[\frac{350}{2800}=\frac{x}{100}\]cross-multiply:\[350*100=2800x\]\[3500=2800x\]you want to solve for \(x\). what number would be the best one to divide both sides by?

OpenStudy (rainbow_rocks03):

2800?

OpenStudy (whpalmer4):

sorry, didn't get enough zeroes.\[35000=2800x\]

OpenStudy (whpalmer4):

yes, divide both sides by 2800. result is?

OpenStudy (rainbow_rocks03):

wait what was I dividing by 2800 on both sides? :(

OpenStudy (rainbow_rocks03):

35000?

OpenStudy (rainbow_rocks03):

Oh, I forgot how to divide some # on both sides of some #. But in this case 2800 is the # I am dividing on both sides of idk which number

OpenStudy (whpalmer4):

anything you do to one side of the equation you must also do to the other side. \[\frac{35000}{2800}=\frac{2800x}{2800}\]\[\frac{35000}{2800}=\frac{\cancel{2800}x}{\cancel{2800}}\]\[x=\frac{35000}{2800}=\]

OpenStudy (rainbow_rocks03):

😞im so close but I cant think what my answer is.

OpenStudy (whpalmer4):

get out a piece of paper and do the long division!

OpenStudy (rainbow_rocks03):

ok 😟

OpenStudy (rainbow_rocks03):

ill try but long division is kind of hard for me

OpenStudy (rainbow_rocks03):

@whpalmer4 Im sorry, I tried. I am trying REALLY hard to find out my answer, but I just cant think anymore. I have been working a lot and I have been working sense 11:00am. But I give up so nevermind. Thx for your help though

OpenStudy (whpalmer4):

then you really need to practice until it is no longer difficult! 2800 | 35000 10*2800 = 28000 which is less than 35000 20*2800 = 56000 which is greater than 35000 answer starts with a 1 35000 - 28000 = 7000 2800 | 7000 2*2800=5600 which is less than 7000 3*2800 = 8400 which is greater, so second digit is 2 7000-5600 = 1400 1400 is 1/2 of 2800 so our quotient is 12 + 1/2 or 12.5